If X has a normal distribution with a mean of 10 and a variance of 4 what is the probability that X will be equal to or greater than 11?
Answer by stanbon(75887)

If X has a normal distribution with a mean of 10 and a variance of 4 what is the probability that X will be equal to or greater than 11?
-----
z(11) = (11-10)/sqrt(4) = 1/2
-----
P(x >= 11) = P(z >= 1/2) = 0.3085
===================
Cheers,
Stan H.
============